### v2

@@ 5 Conclusions and Future Work
We presented FNODE, a framework for data-driven modeling of multibody dynamics that learns accelerations directly rather than states. By reformulating the training objective and leveraging FFT- and finite-difference-based differentiation, FNODE eliminates the adjoint bottleneck of traditional Neural ODEs. Across benchmarks ranging from simple oscillators to chaotic and constrained mechanisms, FNODE trained one to two orders of magnitude faster than MBD-NODE while outperforming black-box baselines such as LSTM and FCNN. Key outcomes include:
  • High accuracy and generalization: FNODE achieved significantly lower MSE than all baselines, performing well in both interpolation and extrapolation.
@@
  • Multiscale robustness: FNODE captured dynamics across widely varying time scales, as shown in the triple-mass-spring system.
  • Reproducibility: An open-source implementation and benchmark suite are provided for future comparisons.
A key ingredient of FNODE is the construction of accurate acceleration targets. FFT-based spectral differentiation offers high precision when trajectories are smooth, but performs poorly at the endpoints of finite, non-periodic data. Finite differences, though lower order, are robust in those cases. By employing these two schemes in a complementary way, FNODE achieves the necessary balance of accuracy and stability, addressing the limitations of using either method alone raissi2019physics .
Two main limitations remain. First, FNODE relies on accurate acceleration estimates from trajectory data; both FFT and FD methods can suffer from noise and, in the case of FFT, Gibbs artifacts. Second, integration error compounds over time, reducing long-term accuracy in chaotic systems such as the double pendulum.
Future work should address the two main limitations noted above. One direction is developing noise-resilient methods for estimating accelerations from sparse or imperfect data. Another is reducing long-term error growth, for example by embedding physical constraints or conservation laws into training or integration. Extending FNODE to higher-dimensional, contact-rich multibody systems would further broaden its impact in engineering applications. Beyond multibody dynamics, applying FNODE ideas to PDE-governed systems, integrating uncertainty quantification, and exploring closed-loop control contexts are promising avenues.
Acknowledgements.
@@ Appendix B Double Pendulum Hamiltonian Formulation
For completeness, we outline the Hamiltonian derivation of the double pendulum dynamics used in Section 3.
The total energy is
$H(q_{1},q_{2},p_{\theta_{1}},p_{\theta_{2}})=T(q_{1},q_{2},p_{\theta_{1}},p_{\theta_{2}})+V(q_{1},q_{2}),$ (30)
@@
$\dot{q}_{i}=\frac{\partial H}{\partial p_{\theta_{i}}},\qquad\dot{p}_{\theta_{i}}=-\frac{\partial H}{\partial q_{i}},$
@@ Appendix C A slider-crank mechanism, with rigid bodies
Based on the settlement illustrated in Section 3.5, we formulates the equation of motion of slider-crank mechanism:
@@
$M_{1}=\begin{bmatrix}m_{1}&0&0\\ 0&m_{1}&0\\ 0&0&I_{1}\end{bmatrix}=\begin{bmatrix}3&0&0\\ 0&3&0\\ 0&0&4\end{bmatrix}$,
$M_{2}=\begin{bmatrix}m_{2}&0&0\\ 0&m_{2}&0\\ 0&0&I_{2}\end{bmatrix}=\begin{bmatrix}6&0&0\\ 0&6&0\\ 0&0&32\end{bmatrix}$,
$M_{3}=\begin{bmatrix}m_{3}&0&0\\ 0&m_{3}&0\\ 0&0&I_{3}\end{bmatrix}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$.
The states of the slider crank mechanism $(x_{1},y_{1},\theta_{1},x_{2},y_{2},\theta_{2},x_{3},y_{3},\theta_{3})$ follows the below constraints $\Phi:\mathbb{R}^{9}\rightarrow\mathbb{R}^{8}$ on the position:
@@
Then the constraints on velocit should be:
@@
$F_{e1}=\begin{bmatrix}0\\ 0\\ \tau-c_{01}\omega_{1}\end{bmatrix}\in\mathbb{R}^{3},$
$F_{e2}=\begin{bmatrix}0\\ -c_{12}(\omega_{1}-\omega_{2})\\ 0\end{bmatrix}\in\mathbb{R}^{3},$
$F_{e3}=\begin{bmatrix}-k\Delta x_{3}-f-c_{23}\omega_{3}-c\dot{x}_{3}\\ 0\\ 0\end{bmatrix}\in\mathbb{R}^{3}.$

### v3

@@ 5 Conclusions and Future Work
We introduced FNODE, a framework for data-driven modeling of multibody dynamics that learns accelerations directly rather than states. By reformulating the training objective and leveraging FFT- and finite-difference-based differentiation, FNODE eliminates the adjoint bottleneck of traditional Neural ODEs. Across benchmarks ranging from simple oscillators to chaotic and constrained mechanisms, FNODE trained one to two orders of magnitude faster than MBD-NODE while outperforming black-box baselines such as LSTM and FCNN. Key outcomes include:
  • Good accuracy and generalization: FNODE achieved lower MSE than all baselines, performing well in both interpolation and extrapolation.
@@
  • Multiscale robustness: FNODE captured dynamics across time scales, as shown in the triple-mass-spring-damper system.
  • Reproducibility: Our open-source implementation and benchmark suite, which includes all NNs and dynamics models used in this study, are provided for reproducibility studies and further research.
FFT-based spectral differentiation yields high accuracy for smooth, uniformly sampled trajectories when periodicity (or approximate periodicity) is a reasonable assumption, since the derivative is obtained from a global Fourier representation. However, for finite, non-periodic trajectories the implicit periodic extension can introduce endpoint artifacts (spectral leakage/Gibbs-type effects), motivating the use of local finite-difference (FD) stencils near boundaries. The FD schemes employed in this work were of lower order (e.g., $\mathcal{O}(\Delta t^{2})$ in the interior and $\mathcal{O}(\Delta t)$ at endpoints for the stencils used here), but are robust for truncated data and do not rely on periodicity. Higher order FD, which have not been tested in this work, could be more accurate but may amplify noise. By employing the FFT and FD schemes in a complementary way, i.e., FFT where its assumptions are approximately satisfied and FD where boundary effects or non-periodicity dominate, we balance accuracy and robustness without relying on either method exclusively.
There are several limitations of the approach which we stopped short of addressing and thus are deferred to future work. First, FNODE relies on accurate acceleration estimates from trajectory data; both FFT and FD methods can suffer from noise and, in the case of FFT, Gibbs artifacts. Further work should be carried out to investigate the sensitivity of the approach to the choice of differentiation scheme. Second, integration error compounds over time, reducing long-term accuracy in chaotic systems such as the double pendulum. However, even when this is the case, the proposed approach still yields accurate predictions for short-term and intermediate-term horizons, a key requirement for controls applications.
Future work should investigate also noise-resilient methods for estimating accelerations from sparse or imperfect data. The goal is to reduce long-term error growth, for example by embedding physical constraints or conservation laws into training or integration, in line with the methodology embraced for the slider-crank example, which has kinematic constraints. An important direction for future work is a systematic numerical study of differentiation-induced errors, including evaluation of the stencil-disagreement indicator $\eta^{\prime}$ and the trajectory reconstruction residual $\eta_{z}$ across datasets and model variants. Finally, one rich direction of research is extending FNODE to problems with friction and contact, in the vein of the slider-crank example. Handling friction is expected to be relatively straightforward for smoothed, penalty-based formulations in which friction is represented through history-dependent ‘bristle’ dynamics, as in models such as LuGre and Dahl. The problem is more challenging for complementarity-based or Lagrange-multiplier formulations, where stick-slip transitions are enforced non-smoothly and can introduce abrupt regime changes. Beyond multibody dynamics, several avenues for future work include applying FNODE ideas to PDE-governed systems, integrating uncertainty quantification, and exploring closed-loop control applications.
The code and scripts used for this work are publicly available as open source for reproducibility studies and further research [55].
Acknowledgements
@@
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used OpenAI ChatGPT in order to improve language and readability of the text. After using this service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Appendix B Double Pendulum Hamiltonian Formulation
For completeness, we outline the Hamiltonian derivation of the double pendulum dynamics used in Section 3. The Hamiltonian is obtained via a Legendre transform of the Lagrangian and is written as
$H(\theta_{1},\theta_{2},p_{\theta_{1}},p_{\theta_{2}})=T(\theta_{1},\theta_{2},\dot{\theta}_{1},\dot{\theta}_{2})+V(\theta_{1},\theta_{2}),$ (37)
@@
The canonical momenta are defined by $p_{\theta_{i}}=\tfrac{\partial T}{\partial\dot{\theta}_{i}}$. Using the expression for $T$ above yields
$p_{\theta_{1}}$ $=(m_{1}+m_{2})l_{1}^{2}\dot{\theta}_{1}+m_{2}l_{1}l_{2}\dot{\theta}_{2}\cos(\theta_{1}-\theta_{2}),$ (40)
$p_{\theta_{2}}$ $=m_{2}l_{2}^{2}\dot{\theta}_{2}+m_{2}l_{1}l_{2}\dot{\theta}_{1}\cos(\theta_{1}-\theta_{2}).$ (41)
Inverting these relations eliminates $\dot{\theta}_{1}$ and $\dot{\theta}_{2}$ in favor of $p_{\theta_{1}}$ and $p_{\theta_{2}}$, producing $H(\theta_{1},\theta_{2},p_{\theta_{1}},p_{\theta_{2}})$ and the explicit Hamiltonian equations below.
@@
$\dot{\theta}_{i}=\frac{\partial H}{\partial p_{\theta_{i}}},\qquad\dot{p}_{\theta_{i}}=-\frac{\partial H}{\partial\theta_{i}},$
@@ Appendix C A slider-crank mechanism, with rigid bodies
Based on the setup illustrated in Section 3.5, we formulate the equation of motion of slider-crank mechanism:
@@
$M_{1}=\begin{bmatrix}m_{1}&0&0\\ 0&m_{1}&0\\ 0&0&I_{1}\end{bmatrix}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0.1\end{bmatrix}$,
$M_{2}=\begin{bmatrix}m_{2}&0&0\\ 0&m_{2}&0\\ 0&0&I_{2}\end{bmatrix}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0.1\end{bmatrix}$,
$M_{3}=\begin{bmatrix}m_{3}&0&0\\ 0&m_{3}&0\\ 0&0&I_{3}\end{bmatrix}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0.1\end{bmatrix}$.
The states of the slider crank mechanism $(x_{1},y_{1},\theta_{1},x_{2},y_{2},\theta_{2},x_{3},y_{3},\theta_{3})$ follow the below constraints $\Phi:\mathbb{R}^{9}\rightarrow\mathbb{R}^{8}$ on the position:
@@
Then the constraints Jacobian assumes the following form:
@@
$F_{e1}=\begin{bmatrix}0\\ 0\\ \tau-c_{01}\dot{\theta}_{1}\end{bmatrix}\in\mathbb{R}^{3},$ (52a)
$F_{e2}=\begin{bmatrix}0\\ -c_{12}(\dot{\theta}_{1}-\dot{\theta}_{2})\\ 0\end{bmatrix}\in\mathbb{R}^{3},$ (52b)
$F_{e3}=\begin{bmatrix}-k\Delta x_{3}-f-\mu|\lambda_{7}|-c\dot{x}_{3}\\ 0\\ 0\end{bmatrix}\in\mathbb{R}^{3}\;,$ (52c)
where $|\lambda_{7}|$ is the absolute value of the scalar Lagrange multiplier (reaction force) associated with the constraint $y_{3}=0$ in Eq. (49), and $\mu$ is the friction coefficient.
@@
Appendix D Analytical Linearization for Cart-Pole
Based on the setup in Section 19, we formulate the cart-pole as a nonlinear continuous system:
$\dot{z}=f(z,u)$ (55)
where $z=[\theta,x,\dot{\theta},\dot{x}]^{T}$ is the system state, $u$ is the control input, $\dot{z}=[\dot{\theta},\dot{x},\ddot{\theta},\ddot{x}]^{T}$ is the time derivative of $z$, and $f$ is the nonlinear function that defines the system dynamics.
To implement the MPC, we linearize the nonlinear system around the equilibrium point $(z_{0},u_{0})$ by first-order Taylor expansion:
$f(z,u)\approx f(z_{0},u_{0})+\frac{\partial f}{\partial z}\bigg|_{(z_{0},u_{0})}\Delta z+\frac{\partial f}{\partial u}\bigg|_{(z_{0},u_{0})}\Delta u$ (56)
In equilibrium state, $z_{0}=0,u_{0}=0,f(z_{0},u_{0})=0$, and the deviations from equilibrium are $\Delta z=z-z_{0}=z$ and $\Delta u=u-u_{0}=u$. The linearized system in Eq. (56) can be expressed as:
$\dot{z}=\frac{\partial f}{\partial z}\bigg|_{(z_{0},u_{0})}z+\frac{\partial f}{\partial u}\bigg|_{(z_{0},u_{0})}u$ (57)
Then we define the linearized system matrices $A$ and $B$, where
$A$ $=\frac{\partial f}{\partial z}\bigg|_{(z_{0},u_{0})}=\begin{bmatrix}\frac{\partial\dot{\theta}}{\partial\theta}&\frac{\partial\dot{\theta}}{\partial x}&\frac{\partial\dot{\theta}}{\partial\dot{\theta}}&\frac{\partial\dot{\theta}}{\partial\dot{x}}\\[8.0pt] \frac{\partial\dot{x}}{\partial\theta}&\frac{\partial\dot{x}}{\partial x}&\frac{\partial\dot{x}}{\partial\dot{\theta}}&\frac{\partial\dot{x}}{\partial\dot{x}}\\[8.0pt] \frac{\partial\ddot{\theta}}{\partial\theta}&\frac{\partial\ddot{\theta}}{\partial x}&\frac{\partial\ddot{\theta}}{\partial\dot{\theta}}&\frac{\partial\ddot{\theta}}{\partial\dot{x}}\\[8.0pt] \frac{\partial\ddot{x}}{\partial\theta}&\frac{\partial\ddot{x}}{\partial x}&\frac{\partial\ddot{x}}{\partial\dot{\theta}}&\frac{\partial\ddot{x}}{\partial\dot{x}}\end{bmatrix}$ (58)
$B$ $=\frac{\partial f}{\partial u}\bigg|_{(z_{0},u_{0})}=\begin{bmatrix}\frac{\partial\dot{\theta}}{\partial u}\\[8.0pt] \frac{\partial\dot{x}}{\partial u}\\[8.0pt] \frac{\partial\ddot{\theta}}{\partial u}\\[8.0pt] \frac{\partial\ddot{x}}{\partial u}\end{bmatrix}$
As it can be assumed that the pole remains near $\theta=0$, we can linearize the system around this equilibrium point to obtain a linearized model for MPC design.
$\sin\theta$ $\approx\theta$ (59)
$\cos\theta$ $\approx 1$
$\dot{\theta}^{2}$ $\approx 0$
Then substitute Eq. (59) into the nonlinear system dynamics Eq. (31) to obtain the linearized model.
$\begin{split}ml^{2}\ddot{\theta}+ml\ddot{x}-mgl\theta&=0\\ ml\ddot{\theta}+(M+m)\ddot{x}&=u\end{split}$ (60)
By rearranging Eq. (60), the accelerations $\ddot{\theta}$ and $\ddot{x}$ can be expressed as:
$\ddot{\theta}$ $=\frac{(M+m)g}{Ml}\theta-\frac{1}{Ml}u$ (61)
$\ddot{x}$ $=-\frac{mg}{M}\theta+\frac{1}{M}u$
Thus, the linearized system matrices $A$ and $B$ can be derived as:
$A=\begin{bmatrix}0&0&1&0\\ 0&0&0&1\\ \frac{(M+m)g}{Ml}&0&0&0\\ -\frac{mg}{M}&0&0&0\end{bmatrix},\quad B=\begin{bmatrix}0\\ 0\\ -\frac{1}{Ml}\\ \frac{1}{M}\end{bmatrix}$ (62)